OpenStudy (anonymous):

Kara's Custom Tees experienced fixed costs of 300 and variable costs of $4 per shirt. Write an equation that can be used to determine the total expenses, find the cost, C, of producing 25 shirts, and graph the equation. Let x equal the total number of shirts.

OpenStudy (anonymous):

m = (5 - -1) / (-10 - -6) = 6 / -4 = -3/2 y = mx + b -1 = (-3/2)(-6) + b -1 = 9 + b b = - 10 Then we have y = -3x/2 - 10 Mulitply both sides by 2 2y = -3x - 20 Add 3x to both sides 3x + 2y = -20 5. C(x) = 4x + 300 Plug 25 in for x C(25) = 4(25) + 300 C(25) = 400 Does that make sense? :)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Hmm. What does not make sense? :)

OpenStudy (anonymous):

So do I divide 25 by 400 for the answer

OpenStudy (anonymous):

No, you would multiply 4 x 25 and add 300 for your answer x(y) means multiply. :)

OpenStudy (anonymous):

100 + 300

OpenStudy (anonymous):

Yep, which equals 400, thus making your answer 400! :)

OpenStudy (anonymous):

400 is not the answer when I type it in

OpenStudy (anonymous):

Hmm.

OpenStudy (anonymous):

@Nnesha can you check my response, not sure what i did wrong.

Nnesha (nnesha):

(5 - -1) / (-10 - -6) where come from ???

OpenStudy (anonymous):

That what I was thinking

Nnesha (nnesha):

per mean what ???

OpenStudy (anonymous):

multiply I think

Nnesha (nnesha):

per also represent slope and slope have x like y = mx+b so 4 per mean 4x

OpenStudy (anonymous):

ok

Nnesha (nnesha):

so y = 4x and your y-intercept 300 so y = 4x + 300

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So the answer is y = 4x + 300

Nnesha (nnesha):

we have to find cost so  x represent what ???

OpenStudy (anonymous):

slope

Nnesha (nnesha):

yep right in the question they told you x represent what ???

Nnesha (nnesha):

Let x equal the total number of shirts.

OpenStudy (anonymous):

Yeah I just read that

Nnesha (nnesha):

so replace x by 25

OpenStudy (anonymous):

100

Nnesha (nnesha):

find the cost, C, of producing 25 shirts, now solve